Carol has received a loan of 50000pesos. She has invested part of it at 10% annual interest and the remaining amount at 14%. If she receives 6400pesos at the end of one year, how much has she invested at each rate?
Answer by checkley77(12844)

.10x+.14(50,000-x)=6,400
.10x+7,000-.14x=6,400
-.04x=6,400-7,000
-.-4x=-600
x=-600/-.04
x=15,000 invested @ 10%.
50,000-15,000=35,000 invested @ 14%.
proof
.10*15,000+.14*35,000=6,400
1,500+ 4,900=6,400
6,400=6,400